# Mining experimental data from Materials Science literature with Large Language Models

### v1

@@ 2.1.2 Formula matching
Matching materials poses challenges with generative models; while encoder and sequence labelling models maintain unchanged outputs from inputs, evaluating generative models can be complex due to potentially divergent yet semantically equivalent outputs. Previous works [32] resort to manual evaluation due to these challenges. Notably, as of the time of writing, no specialised approach tailored for material expressions existed. Utilising Sentence BERT, trained on general text, does not ensure accurate material embeddings, raising concerns about the meaningfulness of final matches. To address issues arising from variable sets and to enhance evaluation precision, we propose a novel method named formula_matching, involving element-by-element pairwise comparisons on normalised formulas for extracted material denominations.
This approach extends strict matching and is activated only when the two input strings differ. In such instances, as depicted in Figure 1, the material expressions slated for comparison undergo normalisation to their formulas using a material parser developed in our prior work [26] (Figure 1 top). The material parser is adept at handling noisy material expressions, striving to parse them effectively. The anticipated output includes a structured representation with the chemical formula presented both as a raw string and as a dictionary, detailing elements and their respective amounts. Subsequently, these structures are compared element by element, as depicted in Figure 1 bottom. The summarised evaluation scores described in Section 3.2 are calculated using the formula matching.
@@ 2.2 Relation Extraction
The baseline is established by a rule-based algorithm from our previous work [26] which was evaluated with SuperMat and for which we report the aggregated result in Section 2.2.
@@
The previous considerations remain applicable to both system and user prompts, with the task description reiterated within both the system and user prompts.
@@
We add specific rules to avoid the creation of invalid groups of relations: responses containing entities that are not supplied in the user prompt, or empty relation blocks.
The prompt for few-shot prompting was assembled by injecting one example with two materials listed between two ”——–” in the zero-shot prompt:
@@
Example:
The researchers of Mg have discovered that MgB2 is superconducting at 29 K at ambient pressure.
entities:
 materials: MgB2, Mg
 tcs: 29K
@@
 tc: 29K,
 pressure: ambient pressure
@@
    - if the material is not specified, ignore the relation block,
    - if tc is not specified in absolute values, ignore the relation block
@@ 2.3 Consideration about the fine-tuning
     - material1
     - material2
     - material3
Using this technique, we could fine-tune a model that was still answering conversationally and then use the GPT-3.5-Turbo base model to transform the response into JSON format.
To fine-tune the model for the RE task, we introduced the sorting variability in the entity lists provided in the prompt (Section 2.2). This approach does not modify the size of the data set and reduces the possibility that the model learns to aggregate entities in the order they appear in the document. This is the default approach we define as ”FT.base” compared to others. In Section 3.3.1 we discuss the impact of two additional strategies for preparing the fine-tuning data. First, ”FT.document_order” keeps the lists of entities as they appear in the document. For example, the made-up sentence ”The two materials MgB2 and MgB3 showed Tc of 39K and 40K, respectively” will lead to two lists of entities ”MgB2, MgB3” and ”39K and 40K” which could be assigned in order (MgB2, 39K) and (MgB3, 40K). Intuitively, this leads to poor performance, as we see when evaluating with shuffling conditions (Section 3.3). The second strategy, ”FT.augmented”, is to augment the size of the dataset generating multiple training records with further shuffled entity list, for each example in ”FT.base”. The data used with this strategy is roughly double that of ”FT.base” (Table 1). We expect this strategy to obtain similar or better results than ”FT.base”.
@@ 3 Results and discussions
In this section, we present and discuss the aggregated results of our evaluations. The completed raw results are available in the Supplementary Material section.
@@ 3.1 NER on properties extraction
The property extraction assessment was performed using the MeasEval dataset, with the baseline established by grobid volumes, achieving an approximately 85% score using a holdout dataset created in conjunction with the application. The evaluation of grobid-quantities [27] against MeasEval yielded a score of around 59%. This disparity was anticipated, given the slightly divergent annotation strategies employed by the MeasEval developers compared to those used in the development of grobid-quantities (e.g., considerations such as approximate values and other proximity expressions were not taken into account).
Unexpectedly, none of the models managed to outperform grobid-quantities in zero-shot prompting, as depicted in Figure 2. This outcome is surprising considering that a) the expression of properties lacks a specific domain constraint (aside from potential variations in frequency distribution), and b) measurements of physical quantities are likely prevalent in the extensive text corpus used to pre-train the OpenAI models.
In the realm of few-shot prompting (Figure 2), a marginal improvement was observed only for GPT-4 and GPT-4-Turbo, resulting in an F1-score gain ranging around 2%. However, this improvement is not significant. We theorise that the hints provided to the LLMs may introduce bias, and when these hints are incorrect or incomplete, the LLMs struggle to guide the generation effectively, impacting the quality of the output results. Significantly, the fine-tuned model (Figure 2) appears to exhibit a slight enhancement compared to zero-shot, few-shot, and the baseline. Interestingly, in this specific instance where both the baseline and fine-tuned models are trained and evaluated on the same data, the LLM demonstrates an approximate 3% increase in the F1-score.
@@ 3.2 NER on materials expressions extraction
In zero-shot prompting (Figure 3), both GPT-4 and GPT-4-Turbo achieved comparable F1-scores, hovering around 50%. Notably, all LLMs scored at least 10% lower than the baseline [26]. This disparity is expected, given that material expressions may involve extensive sequences and encompass multiple pieces of information that are not easily conveyed in the prompt. Few-shot prompting (Figure 3) yielded improved results, with GPT-3.5-Turbo and GPT-4 slightly surpassing the baseline. The introduction of hints in the prompt indeed enhances performance, but, as previously discussed, it appears to strongly influence the LLMs, not able to mitigate the impact of invalid hints that may be provided. Equally unexpected, fine-tuning did not outperform few-shot prompting. This outcome suggests that the additional training did not significantly enhance the LLMs’ ability to handle material expressions.
@@ 3.3 Relation extraction
The evaluation of RE utilised the complete SuperMat dataset, with the results illustrated in Figure 4, providing a comparison of the effects of shuffling across different models.
GPT-3.5-Turbo zero-shot and few-shot prompting demonstrate a significant difference between shuffled and non-shuffled evaluation (Section LABEL:subsub:shuffled-non_shuffled-eval), suggesting a sequential connection of entities without specific contextual reasoning. Notably, the fine-tuned GPT-3.5-Turbo model outperforms the baseline by approximately 15% F1-score and does not show relevant differences when the evaluation is performed under shuffling conditions.
Figure 5 specifically highlights the shuffled version of each model and extraction type. Except for GPT-3.5-Turbo, few-shot prompting shows an improvement compared to zero-shot prompting (Figure 5), achieved by incorporating additional examples in each prompt. GPT-4 and GPT-4-Turbo also exhibit stable results under shuffling conditions, achieving an F1-score of around 15-18% lower than fine-tuned GPT-3.5-Turbo.
@@ 3.3.1 Data variability for fine-tuning
In Section 2.3 we describe two additional ways to prepare the data for fine-tuning. As illustrated in Figure 6, the GPT-3.5-Turbo model fine-tuned with the strategy ”FT.document_order” showed an inability to generalise when evaluated under shuffling conditions, where the model loses around 30% in F1-score. This suggests that adding entropy (for example, by shuffling the data) should be performed as a best practice, which could result in models with larger reasoning capabilities.
When we increased the size of the dataset used in fine-tuning to almost double (Table 1), the resulting model did not show any kind of improvement compared to the FT.base. These results confirm that in fine-tuning, size does not matter, while data variability and quality do.
@@ 5 Conclusion
When it comes to NER, LLMs appear to underperform significantly compared to SLMs in both material and property extraction (Q1) This finding is particularly surprising considering properties since these expressions are not confined to a specific domain.
In material extraction, GPT-3.5-Turbo with fine-tuning failed to outperform the baseline, and the same holds for any model with few-shot prompting. For property extraction, GPT-4 and GPT-4-Turbo with zero-shot prompting perform on par with the baseline. GPT-3.5-Turbo with few-shot and fine-tuning, on the other hand, outperform the baseline by a marginal increase in points. Our results suggest that, for material expressions, small specialised models remain the most accurate choice.
The scenario improves for RE (Q2). With two examples, few-shot prompting demonstrates a significant improvement over the baseline. GPT-4-Turbo exhibits enhanced reasoning capabilities compared to GPT-4 and GPT-3.5-Turbo. GPT-3.5-Turbo performs poorly in both zero-shot and few-shot prompting showing a substantial score decrease when entities are shuffled, aligning with previous observations.
Fine-tuning, nevertheless, yields scores superior to both the baseline and other models, showing results stability when comparing shuffled and not-shuffled evaluation.
In conclusion, to answer Q2, GPT-4 and GPT-4-Turbo showcase effective reasoning capabilities for accurately relating concepts and extracting relations without the need for fine-tuning. However, fine-tuning GPT-3.5-Turbo out yields the best results with a relatively small dataset. GPT-4-Turbo, which costs one third of GPT-4, remains a robust choice given its reasoning capabilities. However, for Q1, for the extraction of complex entities such as materials, we find that training small specialised models remains a more effective approach.
@@ A.3.1 Zero-shot
Table 16: Performance Metrics for the GPT3.5-turbo model on RE in materials-properties extraction. In this run, entities are not shuffled when included in the prompt. P: Precision, R: Recall, F1: harmonic average of P and R, Supp: Support, number of extracted entities.
@@
Run1 Strict matching 80.41 65.7 72.32 934
Soft matching 80.73 65.97 72.6 934
Run2 Strict matching 80.41 54.59 65.03 776
Soft matching 80.8 54.86 65.35 776
Run3 Strict matching 80.25 55.47 65.6 790
Soft matching 80.51 55.64 65.8 790
@@
Strict matching 67.65 4.05 833
Soft matching 67.91 4.06
@@
Run1 Strict matching 76.12 60.8 67.61 913
Soft matching 76.67 61.24 68.09 913
Run2 Strict matching 72.08 48.56 58.02 770
Soft matching 73.12 49.26 58.86 770
Run3 Strict matching 74.9 49.61 59.68 757
Soft matching 75.17 49.78 59.89 757
@@
Strict matching 61.77 5.12 813
Soft matching 62.28 5.05

### v3

@@ 2.1.2 Formula matching
Matching materials poses challenges with generative models, while encoder and sequence labelling models maintain unchanged the output from the input sequences. Therefore, evaluating generative models can be complex due to potentially divergent yet semantically equivalent output sequences. Previous works [34] resort to manual evaluation due to these challenges. Notably, as of the time of writing, no specialised approach tailored for material expressions existed. Utilising Sentence BERT, trained on general text, does not ensure accurate material embeddings, raising concerns about the meaningfulness of final matches. To address issues arising from variable sets and to enhance evaluation precision, we propose a novel method named formula_matching, involving element-by-element pairwise comparisons on normalised formulas for extracted material denominations.
This approach extends strict matching and is activated only when the two input strings differ. In such instances, as depicted in Figure 1, the material expressions slated for comparison undergo normalisation to their formulas using a material parser developed in our prior work [27] (Figure 1 top). The material parser is adept at handling noisy material expressions and strives to parse them effectively. The anticipated output includes a structured representation with the chemical formula presented as a raw string and a dictionary detailing elements and their respective amounts. Subsequently, these structures are compared element by element, as depicted in Figure 1 bottom. The summarised evaluation scores described in Section 3.4 are calculated using the formula matching.
Evaluation and discussion of this method are detailed in Section 3.2.
@@ 2.2 Relation Extraction
The baseline is established by a rule-based algorithm from our previous work [27], which was evaluated with SuperMat and for which we report the aggregated result in Section 2.2.
@@
The previous considerations remain relevant for both system and user prompts, with the task description reiterated in each prompt.
@@
We add specific rules to avoid creating invalid groups of relations and to ignore responses containing entities not supplied in the user prompt or empty relation blocks.
The prompt for few-shot prompting was assembled by injecting three examples listed between the dashed lines (”--------”) in the zero-shot prompt:
@@
Use the following examples separated by "--------" to learn the task:
text 1: The researchers of Mg have discovered that MgB2 and MgB3 are superconducting at 29-31 K at ambient pressure.
entities 1:
materials: MgB2, Mg, MgB3
tcs: 29-31 K
@@
tc: 29-31K,
pressure: ambient pressure:
material: MgB3,
tc: 29-31K,
pressure: ambient pressure:
Text 2: We are studying the material La 3 A 2 Ge 2 (A = Ir, Rh). The critical temperature T C = 4.7 K discovered for La 3 Ir 2 Ge 2 in this work is by about 1.2 K higher than that found for La 3 Rh 2 Ge 2.
entities 2:
materials: La 3 A 2 Ge 2 (A = Ir, Rh), La 3 Ir 2 Ge 2, La 3 Rh 2 Ge 2
tcs: 4.7 K, 1.2 K
Result 2:
material: La 3 Ir 2 Ge 2
tc: 4.7 K
Text 3: The experimental discovery of the high-temperature superconducting state in the compressed hydrogen and sulfur systems H2S (TC = 150 K for p = 150 GPa) and H3S (TC = 203 K for p = 150 GPa)
entities 3:
materials: H2S, H3S
tcs: 150 K, 203 K
pressures: 150 GPa, 150 GPa
Result 3:
material: H2S,
tc: 4.7 K,
pressure: 150 GPa
material: H3S,
tc: 150 K,
pressure: 150 GPa
@@
- if material is not specified, ignore the relation block,
- if tc is not specified in absolute values, ignore the relation block
@@ 2.3 Consideration about the fine-tuning
- material1
- material2
- material3
Using this technique, we could fine-tune a model and shape its behaviour to answer conversationally. Then, we used the GPT-3.5-Turbo base model to transform the response into JSON format.
To fine-tune the model for the RE task, we introduced the sorting variability in the entity lists provided in the prompt (Section 2.2). This approach does not modify the size of the data set and reduces the possibility that the model learns to aggregate entities in the order they appear in the document. This is the default approach we define as ”FT.base” compared to others. In Section 3.5.1, we discuss the impact of two additional strategies for preparing the fine-tuning data. First, ”FT.document_order” keeps the lists of entities as they appear in the document. For example, the made-up sentence ”The two materials MgB2 and MgB3 showed Tc of 39K and 40K, respectively” will lead to two lists of entities ”MgB2, MgB3” and ”39K and 40K” which could be assigned in order (MgB2, 39K) and (MgB3, 40K). Intuitively, this leads to poor performance, as we see when evaluating with shuffling conditions (Section 3.5). The second strategy, ”FT.augmented”, is to augment the size of the dataset, generating multiple training records with a further shuffled entity list for each example in ”FT.base”. The data used with this strategy is roughly double that of ”FT.base” (Table 1). We expect this strategy to obtain similar or better results than ”FT.base”.
@@ 3 Results and discussions
In this section, we present and discuss the formula matching and the aggregated results of our evaluations for the LLMs. The completed raw results are available in the Appendix Figures & Tables.
3.1 Limitation of this study
In this paper, we aim to estimate how well LLMs work in tasks related to materials science. Due to the lack of clean datasets covering the entire materials science domain, we used a dataset that focuses on superconductor material. While our goal is to propose a methodology, we are aware that our results need to be verified empirically in other materials science sub-domains in future works. The following intuitions support our hypothesis: for material NER, we expect that the forms on which materials are presented in other domains would have similar expressions to the ones used in superconductor research, considering that chemical formulas, sample names, and commercial names would unlikely be very different between domains. Furthermore, the properties, expressed as measurement and physical quantities, are common to all domains; although the statistical distribution could be different, we don’t expect dramatic differences within materials science. On the other hand, RE tasks surely require more datasets that focus both on different domains and different flavours of the same task. As an example, the MatSCIRe [35] dataset, which covers battery-related research, proposes a structure that challenges the relation extraction only between two entities (binary extraction) with the addition of the type of relation which could be inferred by the properties being extracted. In conclusion, we will remand the generalisation for further work.
3.2 Formula matching
We evaluated the formula matching to measure two main pieces of information: the gain in the F1-score, and the correctness, as the number of invalid new matches, of the gain. We compared the formula matching with the strict matching because a) it is simple to reproduce and understand visually, and b) the formula matching is built on top of strict matching. We would have more difficulties explaining matches provided by soft matching or SentenceBERT.
We examined the GPT-3.5-Turbo NER extraction (discussed in Section 3.5). 107 out of the 1402 expected records matched correctly using strict matching (P: 22.5%, R: 13.64%, F1: 17.01%). Applying formula matching on the mismatching records, we obtained an additional 176 matches (P: 61.12%, R: 36.00%, F1: 45.31%), for a total gain in F1-score of 28.3 (+266%). For the new 176 records that the formula matching was identifying, we manually examined each pair finding 5 incorrect matches, which corresponds to an error rate of 2.5%.
Most of the mismatches in the strict matching caught up by the formula matching were due to missing adjoined information. The LLMs were not able to include information about doping or shape in the response (e.g. hole-doped La 2-x Sr x CuO 4 was not matching with La 2-x Sr x CuO 4). In other cases, the formula was different by formatting, like: Nd 2-x Ce x CuO 4 and La 2-x Sr x CuO 4. However, the more interesting cases were provided by element or amount substitutions such as: electron-doped infinite-layer superconductors Sr 0.9 La 0.1 Cu 1-x R x O 2 where R = Zn and Ni which was matched Sr0.9La0.1Cu1-xNixO2, or Eu 1-x K x Fe 2 As 2 samples with x = 0.35, 0.45 and 0.5 and Eu 0.5 K 0.5 Fe 2 As 2’. These two cases were particularly complicated to match because they required a deeper understanding of the formula structure.
Among the errors of the formula matching, all of them were provided by the formula which was not correctly parsed, for example in one complicated case with the substrate information: (1-x/2)La 2 O 3 /xSrCO 3 /CuO in molar ratio with x = 0.063, 0.07, 0.09, 0.10, 0.111 and 0.125 which was incorrectly matched with the general La2O3.
@@ 3.3 NER on properties extraction
The property extraction assessment was performed using the MeasEval dataset, with the baseline established by Grobid volumes, achieving an approximately 85% score using a holdout dataset created in conjunction with the application. At the time of writing, the evaluation of grobid-quantities [28] (version 0.7.3666https://github.com/lfoppiano/grobid-quantities/releases/tag/v0.7.3) against MeasEval yielded a score of around 59% F1-score. This disparity was anticipated, given the slightly divergent annotation strategies employed by the MeasEval developers compared to those used in developing grobid-quantities (e.g., considerations such as approximate values and other proximity expressions were not considered).
Unexpectedly, none of the models outperformed grobid-quantities in zero-shot prompting, as depicted in Figure 2. This outcome is surprising considering that a) the expression of properties lacks a specific domain constraint (aside from potential variations in frequency distribution), and b) measurements of physical quantities are likely prevalent in the extensive text corpus used to pre-train the OpenAI models.
In the realm of few-shot prompting (Figure 2), a marginal improvement was observed only for GPT-4 and GPT-4-Turbo, resulting in an F1-score gain ranging around 2%. However, this improvement is not significant. We theorise that the hints provided to the LLMs may introduce bias. When these hints are incorrect or incomplete, the LLMs struggle to guide the generation effectively, impacting the quality of the output results. Significantly, the fine-tuned model (Figure 2) shows a slight enhancement compared to zero-shot, few-shot, and the baseline. Interestingly, in this specific instance where both the baseline and fine-tuned models are trained and evaluated on the same data, the LLM demonstrates an approximate 3% increase in the F1-score.
@@ 3.4 NER on materials expressions extraction
In zero-shot prompting (Figure 3), both GPT-4 and GPT-4-Turbo achieved comparable F1-scores, hovering around 50%. Notably, all LLMs scored at least 10% lower than the baseline [27]. This disparity is expected, given that material expressions may involve extensive sequences and encompass multiple pieces of information not easily conveyed in the prompt. Few-shot prompting (Figure 3) yielded improved results, with GPT-3.5-Turbo and GPT-4 slightly surpassing the baseline. The introduction of hints in the prompt indeed enhances performance, but, as previously discussed, it appears to strongly influence the LLMs, not able to mitigate the impact of invalid hints that may be provided. Equally unexpected, fine-tuning did not outperform few-shot prompting. This outcome suggests that the additional training did not significantly enhance the LLMs’ ability to handle material expressions.
@@ 3.5 Relation extraction
The evaluation of RE utilised the complete SuperMat dataset, with the results illustrated in Figure 4, comparing the effects of shuffling across different models.
GPT-3.5-Turbo zero-shot and few-shot prompting demonstrate a significant difference between shuffled and non-shuffled evaluation (Section 2.2.1), suggesting a sequential connection of entities without specific contextual reasoning. Notably, the fine-tuned GPT-3.5-Turbo model outperforms the baseline by approximately 15% F1-score and does not show relevant differences when the evaluation is performed under shuffling conditions.
Figure 5 specifically highlights the shuffled version of each model and extraction type. Except for GPT-3.5-Turbo, few-shot prompting shows an improvement compared to zero-shot prompting, achieved by incorporating additional examples in each prompt. GPT-4 and GPT-4-Turbo also exhibit stable results under shuffling conditions, achieving an F1-score of around 15-18% lower than fine-tuned GPT-3.5-Turbo.
@@ 3.5.1 Data variability for fine-tuning
In Section 2.3, we describe two additional ways to prepare the data for fine-tuning. As illustrated in Figure 6, the GPT-3.5-Turbo model fine-tuned with the strategy ”FT.document_order” showed an inability to generalise when evaluated under shuffling conditions, where the model loses around 30% in F1-score. This suggests that adding entropy (for example, by shuffling the data) should be performed as a best practice, which could result in models with larger reasoning capabilities.
When we increased the size of the dataset used in fine-tuning to almost double (Table 1), the resulting model did not improve compared to the FT.base. These results confirm that in fine-tuning, size does not matter, while data variability and quality do.
@@ 5 Conclusion
In this study, we have proposed an evaluation framework for estimating how well LLMs perform compared with SLMs and rule-based tasks related to materials science by focusing on sub-domains such as superconductor research. The findings obtained from our work provide initial guidance applicable to other materials science sub-domains in future research.
To evaluate material extraction comparison, we proposed a novel method to parse and match formula elements by elements through an aggregated parser for materials. This new method provides a more realistic F1 score. Compared with strict matching, we obtained a gain in F1-score from 17% to 45% for GPT3.5-Turbo NER at the price of a minimal error rate (2%).
We then evaluated LLMs on two tasks: NER for materials and properties and RE for linking them. LLMs underperform significantly on NER tasks than SLMs in material and property extraction (Q1). This finding is particularly surprising considering properties since these expressions are not confined to a specific domain.
In material extraction, GPT-3.5-Turbo with fine-tuning failed to outperform the baseline, and the same holds for any model with few-shot prompting. For property extraction, GPT-4 and GPT-4-Turbo with zero-shot prompting perform on par with the baseline. GPT-3.5-Turbo with few-shot and fine-tuning, on the other hand, outperforms the baseline by a marginal increase in points. Our results suggest that, for material expressions, small specialised models remain the most accurate choice.
The scenario improves for RE (Q2). With two examples, few-shot prompting demonstrates a significant improvement over the baseline. GPT-4-Turbo exhibits enhanced reasoning capabilities compared to GPT-4 and GPT-3.5-Turbo. GPT-3.5-Turbo performs poorly in both zero-shot and few-shot prompting, showing a substantial score decrease when entities are shuffled, which aligns with previous observations. Nevertheless, fine-tuning yields scores superior to the baseline and other models, showing stability when comparing shuffled and unshuffled evaluations.
In conclusion, to answer Q2, GPT-4 and GPT-4-Turbo showcase effective reasoning capabilities for accurately relating concepts and extracting relations without fine-tuning. However, fine-tuning GPT-3.5-Turbo out yields the best results with a relatively small dataset. GPT-4-Turbo, which costs one-third of GPT-4, remains a robust choice given its reasoning capabilities. However, for Q1, for extracting complex entities such as materials, we find that training small specialised models remains a more effective approach.
@@ A.3.1 Zero-shot
Table 16: Performance Metrics for the GPT3.5-turbo model on RE in materials-properties extraction. In this run, the entities are used as they appear in the input data when included in the prompt. P: Precision, R: Recall, F1: harmonic average of P and R, Supp: Support, number of extracted entities.
@@
Run1 Strict matching 79.90 55.29 65.36 791
Soft matching 80.15 55.47 65.56 791
Run2 Strict matching 80.13 54.33 64.75 775
Soft matching 80.52 54.59 65.07 775
Run3 Strict matching 80.74 55.03 65.45 779
Soft matching 81.00 55.21 65.66 779
@@
Strict matching 65.18 0.38 781
Soft matching 65.43 0.31
Table 17: Performance Metrics for the GPT3.5-turbo model on RE in materials-properties extraction. In this run, we shuffle entities to make it more challenging for the LLM. P: Precision, R: Recall, F1: harmonic average of P and R, Supp: Support, number of extracted entities.
@@
Run1 Strict matching 76.01 51.27 61.23 771
Soft matching 76.91 51.88 61.96 771
Run2 Strict matching 75.82 50.74 60.80 765
Soft matching 76.34 51.09 61.22 765
Run3 Strict matching 76.55 52.84 62.53 789
Soft matching 77.19 53.28 63.04 789
@@
Strict matching 61.52 0.90 775
Soft matching 62.07 0.91